Find the volume of a spherical sector of a sphere whose surface area is 437π m^2, if the height of the zone base is 5 m
Answer by Alex.33(110)

top surface area=25π m^2
top surface circumference=10π m
side surface area=437π-2*25π=287π m^2
height=287π/10π=28.7 m
volume=25π*28.7=2254.092729 m^3
And we're done!
Tip: First figure out what we need(height), then design the approach to it(from the side surface area), and do.